1. A 9.50 cm radius platter starts from the rest. It then uniformly accelerates at 2.25 rad/s^2 for 0.950 s. What is the angular speed of the platter?

2. The siren of a firetruck has a frequency of 1.20 kHz. When the firetruck is moving at 24.6 m/s east, what frequency is heard by someone in a car traveling 4.47 m/s east away from the fire truck. The temperature is 30.0 C
Answer #1

O - wo, W, X and 1) As platter starts from the rest, so initial angular speed speed of blatter = W. = o Angular acceleratim oجا = C + کا محو 2.375 = رها ( اتي إمحو 1375 .3 کا محو زوفه - ) ا 6 So, angular speed of platter is 2014 read is